OpenStudy (anonymous):

An isosceles trapezoid has base angles equal to 45 and bases of lengths 6 and 12. Find the area of the trapezoid.

OpenStudy (anonymous):

is it 54

OpenStudy (anonymous):

first calculate the height get this from right anglad triangle formed from perpendicular line from top corner tan 45 = h / 3 then use th area formula

OpenStudy (anonymous):

height = 3

OpenStudy (anonymous):

area = 1/2 * 3 * ( 6 + 12) = 27